# EaDeblur-GS: Event assisted 3D Deblur Reconstruction with Gaussian Splatting

### v2

@@ 2 Method
As illustrated in Fig.1, our method begins with the input of blurry RGB images and their corresponding event streams. We first use the Event Double Integral (EDI) technique to generate a set of latent sharp images. These images are then processed with COLMAP, enhancing the initial reconstruction and providing relative precise camera pose estimation. From this reconstruction, we create a set of 3D Gaussians. The positions of these Gaussians, along with the estimated camera poses, are input into our Adaptive Deviation Estimator (ADE) network to determine positional deviations, which are added to the original Gaussian centers. The adjusted 3D Gaussians are projected onto each viewpoint, including corresponding latent viewpoints, to render sharp images. We also introduce a Blurriness Loss to simulate realistic blurriness and an Event Integration Loss to improve object shape accuracy in the Gaussian model. This process enables the model to learn precise 3D volume representations and achieve superior 3D reconstructions. The overall method is depicted in Fig.1.
Figure 1: The overview of EaDeblur-GS
@@ 3.1 Comparisons and Results
The quantitative comparison results are presented in Table 1 and the qualitative comparison is also provided in supplementary materials.
@@ 3.2 Ablation studies
Blurriness loss and Event Integration Loss. To validate the effectiveness of training losses, we conduct the experiments on the "hotdog" scene from E2NeRF dataset. The results in the Tab.2 demonstrates that both proposed losses significantly improve the deblurring performance, with the event integration loss further slightly enhancing sharp rendering quality.

### v3

@@ 2 Method
As illustrated in Fig.1, our method begins with the input of blurry RGB images and their corresponding event streams. We first use the Event Double Integral (EDI) technique to generate a set of latent sharp images. These images are then processed with COLMAP, enhancing the initial reconstruction and providing relative precise camera pose estimation. From this reconstruction, we create a set of 3D Gaussians. These Gaussian positions, combined with the estimated camera poses, are then fed into our Adaptive Deviation Estimator (ADE) network to adjust the Gaussian centers by calculating positional deviations. The adjusted 3D Gaussians are projected onto each viewpoint, including corresponding latent viewpoints, to render sharp images. We also introduce a Blurriness Loss to simulate realistic blurriness and an Event Integration Loss to improve object shape accuracy in the Gaussian model. This process enables the model to learn precise 3D volume representations and achieve superior 3D reconstructions. The overall method is depicted in Fig.1.
Figure 1: The Overview of our approach. Our approach integrates blurred RGB images with event data to enhance image clarity. The EDI technique is applied to produce latent sharp images, which are then refined through COLMAP for accurate SFM reconstruction and 3D Gaussian modeling. The ADE module estimates positional deviations based on initial Gaussian positions and camera extrinsics, simulating camera motion. These deviated Gaussians are rendered into multiple views and compute blurriness and event integration losses, facilitating the learning of detailed 3D representations for enhanced reconstruction quality.
@@ 3.1 Comparisons and Results
The quantitative comparison results are presented in Table 1.
@@
Figure 2: Qualitative results on E2NeRF synthetic dataset.
We also assess the qualitative performance of our method compared to other approaches. As illustrated in Fig.2, the original 3D Gaussian Splatting struggles to produce clear 3D representations in certain scenes, even with EDI-preprocessed relatively sharp input images. While the Deblurring 3D Gaussian Splatting method can reconstruct relatively sharp objects, it still fails to capture fine details in some cases, such as with the “materials” object. In contrast, our method demonstrates the capability to reconstruct fine-grained, clear objects with greater accuracy.
@@ 3.2 Ablation studies
Blurriness loss and Event Integration Loss. To validate the effectiveness of training losses, we conduct the experiments on the “hotdog" scene from E2NeRF dataset. The results in the Tab.2 demonstrates that both proposed losses significantly improve the deblurring performance, with the event integration loss further slightly enhancing sharp rendering quality.
@@ 4 Conclusion
Acknowledgments
This work was supported by the National Science and Technology Major Project under Grant 2020AAA0107300, and the Key Research and Development Plan Project of Xinjiang Uygur Autonomous Region under Grant 2023B01006-1.